Could someone please help me with this problem. I don't know how to find the perimeter.
The area of a square is 25 square meters. What is its perimeter?
Thanks, Jessie.

Answer by checkley71(8403)

THE AREA OF A SQUARE=SIDE*SIDE
25=S^2
--------------------------------------------------------
SORRY: THIS NEXT LINE WAS S62 & SHOULD HAVE BEEN S^2
--------------------------------------------------------
S^2=25
S=5 METERS IS THE LENGTH OF THE SIDE OF THUIS SQUARE.
PERIMETER OF A SQUARE=4*S
PERIMETER=4*5=20 METERS